OpenStudy (anonymous):

Simplify the following expression. 2x - 2y + 5z - 2x - y + 3z x - 3y + 8z 2x - 2y + 8z 3y + 8z -3y + 8z

OpenStudy (anonymous):

2x - 2y + 5z - 2x - y + 3z combine like terms

OpenStudy (anonymous):

2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

4th option

OpenStudy (anonymous):

1st option @hayleystonem

OpenStudy (anonymous):

explain why @rvgupta

OpenStudy (anonymous):

@katragaddasaichandra you should also explain please don't just give answers.

OpenStudy (anonymous):

i can explain to u @rvgupta

OpenStudy (anonymous):

i just solve it in my notebook and telling answers i am not guessing if u want answer i can type the whole thing. this is the answer : switch like terms next to each other (2x - 2y + 5z - 2x - y + 3z) change to ( 2x - 2x + y - 2y - 5z + 3z) next the y thats all alone add one in front of it.. thats just saying that theres only one y add/subtract like terms 2x - 2x = 0, 1y - 2y = 1y, 5z + 3z = 8z so you'll have ( 0 + 1y - 8z) leave the zero and youll answer : 1y - 8z

OpenStudy (anonymous):

answer is 1y-8z

OpenStudy (anonymous):

how @katragaddasaichandra

OpenStudy (anonymous):

@katragaddasaichandra Thanks, it's just that this site is for help :) not just answers :)

OpenStudy (anonymous):

so plz press best response

OpenStudy (anonymous):

2x-2x-2y-y+5z+3z 0 – 3y +8z

OpenStudy (anonymous):

@katragaddasaichandra its -2y-y = -3y

OpenStudy (anonymous):

no u r wrong i just explained u @surdawi

OpenStudy (anonymous):

switch like terms next to each other (2x - 2y + 5z - 2x - y + 3z) change to ( 2x - 2x + y - 2y - 5z + 3z) next the y thats all alone add one in front of it.. thats just saying that theres only one y add/subtract like terms 2x - 2x = 0, 1y - 2y = 1y, 5z + 3z = 8z so you'll have ( 0 + 1y - 8z) leave the zero and youll answer : 1y - 8z

OpenStudy (anonymous):

where did you get the +y from?

OpenStudy (anonymous):

1st 1y - 2y = 1y is incorrect to begin with; correction: 1y - 2y = -1y however the correct one is -1y - 2y = -3y all 3 people how gave you medal dont know math at all

OpenStudy (anonymous):

@katragaddasaichandra can you just look at the problem again and to your answer you said (2x - 2y + 5z - 2x - y + 3z) change to ( 2x - 2x + y - 2y - 5z + 3z) 1. why do we have -y change to +y and +5z to -5z? 2. 2x - 2x = 0, 1y - 2y = 1y, 5z + 3z = 8z 1y - 2y = -1y; you said step earlier its -5z+3z so that should be -2z 3. ( 0 + 1y - 8z) in step two you said it was +8z, how did it change to negaive 3.